With 3% moisture, 2.5 kg of steam has an enthalpy of 6600 kJ. It is heated at constant pressure to a final condition of 90
1
Expert's answer
2021-08-05T03:19:26-0400

"3 \%" of moisture, "x= 1-0.03=0.97 \\
h= \frac{H}{m}=\frac{6600}{2.5}=2640 kJ/kg\\
At \space x = 0.97, h= 2640 kJ/kg; P=0.2MPa\\"

From the steam table

"P= 0.2 MPa\\
T_{sat}=120.21^0C\\
h_f= 504.7 kJ/kg\\
h_{fg}=2201.5 kJ/kg\\
h= h_f+ xh_{fg}= 504.7+0.97*2201.5 =2640.155 kJ/kg\\"

Degree of super heat "=T_2-T_{sat}"

"80^0C=T_2 -120.21 ^0C \implies T_2 = 200.21^0C"

"At \space P= 0.2MPa\\
T= 200.12 ^0C\\
h= 2871.2 kJ/kg\\
H_2= m*h_2= 2.5*2871.2 = 7178 kJ\\"

Heat received is given as H2-H1

"7178-6600=578 kJ"